# GT2-GS: Geometry-aware Texture Transfer for Gaussian Splatting

### v1

@@ 4.3. Geometry-consistent Texture Loss
(7) $F_{t}(i,j)=\mathop{\arg\min}\limits_{f_{k,\theta},~{}k=k_{(i,j)}}dist(F_{r}(i,j),f_{k,\theta})+\lambda_{p}X(i,j)|\theta^{\prime}+\beta-\theta|,$
@@
where $\lambda_{p}$ is the prior texture orientation control coefficient and $k_{i,j}$ represents the depth group corresponding to the feature position. Subsequently, texture transfer can be achieved by minimizing the cosine similarity between the feature vectors at corresponding positions in the rendered feature map $F_{r}$ and the target feature map $F_{t}$. The geometric-consistent texture transfer loss function is formulated as follows:
(8) $L_{gt}=\frac{1}{N}\sum_{i,j}dist(F_{r}(i,j),F_{t}(i,j)).$
Using only our GT Loss makes it difficult to preserve the original scene content. Therefore, we introduce additional loss functions during the texture transfer stage to retain the scene’s structural information. Needle-like Gaussians negatively affect the quality of texture transfer, making the content difficult to recognize. We measure Gaussian elongation degree using the ratio $l_{max}/l_{mid}$ between the longest axis length $l_{max}$ and the median axis length $l_{mid}$. The regularization term for suppressing needle-like Gaussians is formulated as
(9) $L_{reg}=\frac{1}{N_{G}}\sum_{i}\frac{l_{max}^{i}}{l_{mid}^{i}},$
@@
where $N_{G}$ is the number of Gaussians. The total loss during the texture transfer phase is expressed as
(10) $L_{tot}={\lambda_{gt}}L_{gt}+{\lambda_{c}}L_{content}+{\lambda_{tv}}L_{tv}+{\lambda_{reg}L_{reg}},$
@@
where $L_{tv}$ is the total variation loss, $\lambda_{gt},\lambda_{c},\lambda_{tv},\lambda_{reg}$ are the coefficients of the corresponding loss functions.
Figure 4. Qualitative Comparison of Texture Transfer.
Figure 5. Qualitative Comparison of Style Transfer.
4.4. Geometry Preservation Strategy
@@
3DGS is an explicit representation, where the scene’s ability to learn textures is closely related to the distribution of Gaussians. In low-texture regions, the density of Gaussians is typically lower. While these regions achieve good rendering quality in the original rendering process, they struggle to learn new texture appearances. The Gaussian densification strategy increases the number of Gaussians in low-texture regions during the texture migration phase, enabling the learning of complex textures. However, Gaussians are densified solely based on gradients. Since texture transfer lacks ground truth during optimization, the densification strategy may introduce erroneous Gaussians floating in space.To address this issue, we propose a two-stage iterative optimization strategy.
Specifically, We perform $N_{iter}$ iterations of optimization. In each iteration, our texture transfer framework first performs the texture transfer stage, followed by the geometry correction stage. During the texture transfer stage, the Gaussian color parameters are set to $c$. We optimize the Gaussian parameters using the total loss function $L_{tot}$ defined in Eq. 10. For the GT Loss $L_{gt}$, we dynamically control the injection of prior information. During the iterative optimization process, prior information $\theta^{\prime}$ and $\beta$ in Eq. 7 are introduced in odd-numbered iterations, while it is omitted in even-numbered iterations. This strategy prevents overly strong priors from adversely affecting fine-grained geometric structures in the scene, such as staircases and railings.
Subsequently, we perform the geometry correction stage to address geometric errors introduced during the densification process. At this stage, the Gaussian parameters are set to the additionally embedded color parameters $c^{g}$. This optimization process employs the original 3DGS reconstruction loss function,
(11) $\mathcal{L}_{rec}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L}_{D-SSIM},$
@@
where $\mathcal{L}_{1}$ and $\mathcal{L}_{D-SSIM}$ are calculated between the rendered images and the content images. Through an optimization process with ground truth, the Gaussians in the scene are moved to their correct geometric positions. Through multiple iterations of the aforementioned two-stage optimization process, we achieve a balance between texture transfer and geometry preservation.
Figure 6. Qualitative Experiments on Texture Orientation Control.
5. Experiment
@@
Datasets. For the scene datasets, we utilize the LLFF dataset  (Mildenhall et al., 2019) and the T&T dataset (Knapitsch et al., 2017), which are collected from the real world. Additionally, we use images from the ARF (Zhang et al., 2022) style dataset and the DTD dataset (Cimpoi et al., 2014) as reference image datasets.
@@
Baseline. We compare our method with the state-of-the-art 3D stylization method, including ARF (Zhang et al., 2022), ABC-GS (Liu et al., 2025) and StyleGaussian (Liu et al., 2024). Specifically, ARF is based on NeRF, and ABC-GS and StyleGaussian are based on 3DGS. Optimization-based methods include ARF and ABC-GS, while StyleGaussian is a feed-forward-based method.
5.1. Implementation Details
@@
We perform view-consistent color transfer (Zhang et al., 2022) on the rendered images before and after texture transfer and use them as content images to optimize the Gaussian parameters. For the VGG  (Simonyan and Zisserman, 2014) feature extractor, we employ the conv3 block of VGG-16. For texture feature, the GTA module defaults to a depth group count $K$ to 4, and the rotation angle set $\Phi$ covers 360 degrees. Following the GPS, we set the number of iterations $N_{iter}$ to 10 for texture transfer. During the texture transfer stage, we set $\{\lambda_{gt},\lambda_{c},\lambda_{tv},\lambda_{reg}\}$ = {2, 0.005, 0.02, 0.1}. Some hyperparameters for forward-facing scenes and 360-degree outdoor scenes differ, and we will provide a more detailed explanation in the Appendix. To ensure a fair comparison with other methods, we use the texture direction of the texture image itself as the prior direction during training. All our experiments are conducted on a single NVIDIA RTX 4090 GPU.
5.2. Qualitative Evaluation
@@
To comprehensively evaluate the effectiveness of our method, the qualitative experiments are divided into three parts: texture transfer, style transfer, and texture orientation control. The results are shown in Fig. 4, Fig. 5 and 6, with more qualitative examples provided in the Appendix.
@@
Texture Transfer. Fig. 4 shows the texture transfer results of our method compared with other methods. Visually, our results exhibit higher fidelity to the reference texture and better alignment with human visual perception. For example, in the geometrically complex orchids scene (Row 1 in Fig. 4), our method produces transfer results that appear as if the texture is naturally wrapped around the surface of the orchids. StyleGaussian, as a zero-shot style transfer method, struggles to handle such complex texture transfers. The results of ABC-GS and ARF appear blurry and exhibit texture discontinuities. These methods show suboptimal texture style matching and visual quality. This is primarily due to the use of identical texture features across multiple views during optimization, which contradicts the geometric consistency of the 3D scene. In comparison, our method leverages the GTA module to extract geometry-aware texture features and applies GT Loss for geometry-consistent appearance optimization.
Style Transfer. To validate the generalizability of our method, we selected artistic style images as reference inputs and conducted a qualitative comparison of style transfer with other methods. As shown in Fig. 5, our approach is more effective in transferring the texture elements of the style image to the 3D scene. In contrast, the results of other methods rely more on rendered views and fail to account for the underlying scene geometry. For instance, in the fortress scene shown in the first row of Figure X, ABC-GS and ARF tend to preserve the color difference between tabletop objects and the table during style transfer. In comparison, our method takes full geometric information into account, allowing textures to be accurately transferred onto object surfaces.
@@
Table 1. Quantitative Experiment on Multi-view Consistency.
Methods Short-term Consistency Long-term Consistency
LPIPS($\downarrow$) RMSE($\downarrow$) LPIPS($\downarrow$) RMSE($\downarrow$)
StyleGaussian 0.052 0.051 0.137 0.125
ARF 0.078 0.067 0.175 0.153
ABC-GS 0.051 0.045 0.127 0.118
Ours 0.052 0.045 0.128 0.118
Table 2. Quantitative Experiment on Content Protection.
Ours ABC-GS ARF StyleGaussian
SSIM($\uparrow$) 0.47 0.55 0.34 0.40
CLIP-score($\uparrow$) 0.45 0.39 0.38 0.35
5.3. Quantitative Evaluation
In the 3D scene appearance editing task, maintaining multi-view consistency and preserving scene content information are crucial. To this end, we conducted extensive quantitative experiments from both aspects. We randomly selected 50 pairs of scene-reference image pairs from the scene and reference image datasets to quantitatively evaluate our proposed method.
Multi-view Consistency. We evaluate multi-view consistency using both short-term and long-term consistency metrics. These metrics are computed by warping the rendered image from one viewpoint to another using optical flow (Teed and Deng, 2020). The RMSE and LPIPS scores are then calculated between the warped image and the target view. The results are shown in Table. 1. It can be observed that both our method and ABC-GS achieve high-quality multi-view consistency. Notably, ABC-GS disables the Gaussian densification strategy during optimization. In contrast, our method maintains multi-view consistency even after applying the densification strategy, demonstrating the effectiveness of the proposed GPS.
Content Preservation. For 3D texture transfer, it is essential to ensure that the original scene content remains recognizable while editing the scene’s appearance. We use SSIM (Wang et al., 2004) and CLIP-score (Radford et al., 2021) to assess the preservation of content information. Specifically, SSIM evaluates the structural and informational similarity between two images at the pixel level. In contrast, CLIP-score measures the semantic similarity by computing the cosine similarity between the CLIP embeddings of the two images. As shown in Table. 4, our proposed method achieves significantly higher scores on both evaluation metrics. In particular, our method outperforms previous approaches in terms of the CLIP-score. It indicates that incorporating geometric information enables accurate texture transfer to the scene appearance while preserving the semantic content of the scene.
Table 3. The results of user study.
Ours ABC-GS ARF StyleGaussian
Texture Alignment($\uparrow$) 4.23 3.02 3.08 2.15
Visual Quality($\uparrow$) 3.90 3.37 3.18 2.22
Figure 7. Ablation study for GT Loss.
Figure 8. Ablation study for GTA and GPS.
5.4. User Study
Given the highly subjective nature of texture transfer evaluation, it is challenging to rely solely on objective quantitative metrics for a comprehensive assessment. Therefore, we conducted a user study to evaluate the effectiveness of the proposed method in terms of texture transfer quality. We collected X responses from an online questionnaire distributed through social media platforms. The questionnaire consisted of 20 groups of images along with corresponding prompts. Each group included a texture image, an original content image from a random viewpoint, and a rendered image produced by one of the methods under comparison from the same viewpoint. Each rendered image was evaluated in terms of texture alignment and visual quality, with scores ranging from 1 (lowest) to 5 (highest). As shown in Table. 3, our method consistently outperforms the baseline approaches. Specifically, by comparing the differences across various evaluation dimension, it is evident that users showed a clear preference for our proposed method. This indicates that our approach achieves texture transfer results that better align with human visual perception.
5.5. Ablation Study
Impact of GT Loss. GT Loss is proposed to control the texture orientation. In the ablation study, we perform central rotation on the same texture image to generate multiple inputs with different rotation angles. As shown in Fig. 7, removing the GT Loss leads to identical texture transfer results across different input angles. The reason for this phenomenon is that texture images with different rotation angles produce identical texture feature sets after being processed by the GTA module. Without the constraint of GT Loss, the scene directly learns appearance from the aggregated texture feature set, ignoring orientation variations.
Impact of GTA. As shown in Fig. 8, removing the GTA module results in noticeable texture discontinuities and blurring in the texture transfer outputs. This degradation is attributed to the absence of geometry-aware texture features. Without the GTA module, a large portion of texture features aligned with the scene’s geometry are eliminated. The remaining texture features are insufficient to support learning across the full range of geometric structures within the scene. Learning texture features of the same orientation across different viewpoints fails to achieve continuous texture transfer results between views.
Impact of GPS. The 3D texture transfer tasks inherently lack ground truth supervision, which can introduce incorrect geometry during the scene appearance editing process. This issue is particularly pronounced in the 3DGS-based methods, where appearance and geometry are tightly entangled. To address this, we propose GPS, which leverages the original content images to optimize the geometric parameters of Gaussians and correct inaccurate geometry. As illustrated in Fig. 8, removing GPS results in noticeable artifacts across the scene, especially in challenging 360-degree outdoor environments. This highlights the critical role of GPS in maintaining geometric consistency. Through GPS, our texture transfer framework achieves a balance between learning texture features and preserving scene geometry.
6. Conclusion
In this paper, we introduced GT2-GS, a novel Geometry-aware Texture Transfer framework for Gaussian Splatting that achieves high-quality texture transfer results. Unlike previous 3d stylization methods, our approach explicitly considered the intrinsic relationship between geometry and texture. We observed that feature vectors directly extracted from texture images are insufficient to capture the rich geometric information present in 3D scenes. To address this, we proposed the GTA module, which enhanced the texture feature set by incorporating geometric cues such as perspective geometry, edges, and shapes. Furthermore, we designed a GT Loss that leveraged 3D geometric information across viewpoints to enable controllable texture transfer. To preserve accurate geometry during texture learning, we introduced the GPS strategy to maintain structural integrity throughout the optimization process. Extensive quantitative and qualitative experiments demonstrated the effectiveness of our proposed method. Moreover, our framework was capable of generating high-quality stylization results, showcasing its generalizability.
Acknowledgements.
This work is supported by the National Natural Science Foundation of China (62472178, 62376244), and Shanghai Urban Digital Transformation Special Fund Project (202301027). We also thank the financial support from Shanghai China Fortune Co., Ltd.
Appendix A ADDITIONAL ANALYSIS
Quantitative Ablation Study on GTA and GPS. We perform a quantitative ablation study on the Geometry-aware Texture Augmentation (GTA) module and the Geometry Preservation Strategy (GPS). The impact of texture transfer on the original scene content is assessed using SSIM (Wang et al., 2004) and CLIP-score (Radford et al., 2021). We randomly selected 30 pairs of scenes and texture images for the experiment. As shown in Table  4, removing either of these components adversely affects the geometric structure of the scene. Consequently, the scene content becomes less distinguishable, leading to a decrease in semantic similarity with the original scene.
Component-level Impact Analysis of GTA. We conduct a component level ablation study on the proposed GTA module, which includes multi-scale texture enhancement to simulate perspective geometry and multi-angle enhancement to match geometric cues such as edges and contours in the rendered images. As shown in Fig. 9, the multi-angle enhancement significantly impacts the final texture transfer results, while removing the multi-scale enhancement leads to a lack of depth perception in the scene.
Component-level Impact Analysis of GPS. GPS consists of two key components: dynamic prior injection and the geometry correction stage. We conduct ablation studies to investigate their impact on scene geometry. As shown in Fig. 10, removing dynamic prior injection disrupts the geometric structure of textures that are difficult to learn in the specified orientation, resulting in blurry renderings. On the other hand, removing the geometry refinement stage leads to severe Gaussian floating artifacts throughout the scene, severely degrading the 3D geometric integrity.
Impact of Color Transfer. We follow ARF (Zhang et al., 2022) to perform view-consistent color transfer. As shown in Fig. 12, without color transfer, the scene struggles to learn colors consistent with the texture image, which significantly degrades the visual quality of 3D texture transfer.
Appendix B MORE IMPLEMENTATION DETAILS
For the forward-facing scene dataset LLFF and the 360-degree scene dataset T&T, we adopt different experimental settings. In the color transfer stage, we perform 400 and 1000 optimization steps for LLFF and T&T scenes, respectively. During the subsequent texture transfer stage, for the LLFF dataset, we set $\{\lambda_{gt},\lambda_{c},\lambda_{tv},\lambda_{reg}\}$ = {2, 0.005, 0.02, 0.1} and perform $N_{iter}=10$ rounds of GPS optimization, where each round consists of 80 texture transfer steps followed by 400 geometry correction steps. For the T&T dataset, we reduce parameter $\lambda_{c}$ to 0.0005 and use $N_{iter}=6$ rounds with 400 texture transfer steps and 1000 geometry correction steps per round to accommodate more viewpoints.
Figure 9. Component-level Ablation Study of GTA.
Table 4. Quantitative Experiment on Content Protection.
Ours w/o GTA w/o GPS
SSIM($\uparrow$) 0.43 0.37 0.39
CLIP-score($\uparrow$) 0.36 0.35 0.34
Figure 10. Component-level Ablation Study of GPS.
Figure 11. Impact of Content Loss.
Figure 12. Impact of Color Transfer.
Figure 13. Qualitative Comparison of Texture Transfer on the LLFF Dataset.
Figure 14. Qualitative Comparison of Style Transfer on the LLFF Dataset.
Figure 15. Qualitative Comparison of Texture Transfer on the T&T Dataset.
Figure 16. Qualitative Comparison of Style Transfer on the T&T Dataset.
Appendix C MORE QUALITATIVE EVALUATION
@@
For more qualitative evaluation, we compare our method with ARF (Zhang et al., 2022), ABC-GS (Liu et al., 2025) and StyleGaussian (Liu et al., 2024). Fig. 13, Fig. 14, Fig. 15, and Fig. 16 present the qualitative results.
Appendix D DISCUSSION, LIMITATION, AND FUTURE WORK
@@
Our method primarily consists of three key components: Geometry-aware Texture Augmentation (GTA), Geometric-consistent Texture (GT) Loss, and Geometry Preservation Strategy (GPS). Among them, the proposed GTA module and GT Loss leverage 2D and 3D geometric information respectively, and their computational process is independent of specific 3D representations. Therefore, they can be seamlessly applied to other representations, such as NeRF, TensoRF, and Plenoxel. Meanwhile, GPS provides a novel perspective for geometry-aware optimization in 3DGS-based appearance editing tasks without ground truth supervision. However, from another perspective, the upper bound of geometric accuracy after texture transfer in our framework depends on the quality of geometry obtained from the original 3DGS optimization, which may be suboptimal. Due to the introduction of a geometry correction stage, our method introduces additional optimization time. Nevertheless, for forward-facing scenes, our method is able to achieve high-quality texture transfer results within 5 minutes.

### v3

@@ Introduction
$F_{t}(i,j)=\mathop{\arg\min}\limits_{f_{k,\theta}}dist(F_{r}(i,j),f_{k,\theta})+\lambda_{p}|\theta^{\prime}+\beta-\theta|,$ (6)
@@
where $\lambda_{p}$ is the prior texture orientation control coefficient. Subsequently, texture transfer can be achieved by minimizing the cosine similarity between the feature vectors at corresponding positions in the rendered feature map $F_{r}^{v}$ and the target feature map $F_{t}^{v}$ for the current viewpoint $v$. The geometry-aware texture transfer loss function is formulated as follows:
$L_{gt}=\frac{1}{N}\sum_{i,j}dist(F_{r}^{v}(i,j),F_{t}^{v}(i,j)).$ (7)
Adaptive Fine-grained Control Module
Most existing optimization-based style transfer methods rely on VGG features. However, feature maps extracted through multiple layers of convolutional neural networks exhibit significantly lower granularity compared to the original image pixels. When optimizing the scene using only texture transfer losses, this can lead to the following issues: First, due to perspective projection, regions with greater depth tend to concentrate more scene information. Learning coarse-grained texture features may overwrite these important details. Second, scenes often contain geometrically fine-grained structures such as stairs and railings, which can be degraded or lost under coarse texture representations. To address these challenges, we propose an Adaptive Fine-Grained Control Module.
Specifically, we first obtain the depth map $I_{d}$ and frequency density map $I_{f}$ for each viewpoint from the original Gaussian scene and content image, respectively. Both are rescaled to match the spatial dimensions of the feature maps. Furthermore, we introduce a geometric distortion map $\Phi$, defined as the discrepancy in geometric information between the learned texture features and rendered features. The geometric distortion map is computed by measuring the angular difference between the features obtained using Eq. 6 and those derived without any prior information. The adaptive fine-grained control map is formulated as
$W^{v}=\lambda_{d}(1-I_{d}^{v})+\lambda_{f}(1-I_{f}^{v})+\lambda_{\Phi}(1-\Phi),$ (8)
where $I_{d}^{v}$, $I_{f}^{v}$ and $\Phi$ are all normalized, and $\lambda_{d}$, $\lambda_{f}$ and $\lambda_{\Phi}$ are weight coefficients. In most cases, we aim to alter the appearance of foreground objects while simultaneously satisfying requirements for both shallow depth and high frequency. We therefore employ an additive formulation to balance the relationship between depth and frequency. Furthermore, we regularize the learning process to favor textures with low distortion, thereby better preserving geometric fidelity during texture adaptation.
The derived adaptive weight matrix $W^{v}$ is subsequently applied to Eq. 7 as
$L_{wgt}=\frac{1}{N}\sum_{i,j}W^{v}(i,j)\ dist(F_{r}^{v}(i,j),F_{t}^{v}(i,j)),$ (9)
@@
where $L_{wgt}$ is weighted GT2 Loss. The total loss during the texture transfer phase is expressed as
$L_{tot}={\lambda_{wgt}}L_{wgt}+{\lambda_{c}}L_{content}+{\lambda_{tv}}L_{tv},$ (10)
@@
where $L_{tv}$ is the total variation loss, $\lambda_{wgt}$, $\lambda_{c}$, $\lambda_{tv}$ are the coefficients of the corresponding loss functions.
Figure 4: Comparison of Geometry Preservation Results. It can be observed that as the number of Gaussians increases, depth regularization struggles to preserve the geometric information of the scene. $\lambda_{dep}$ is the weighting coefficient for the depth regularization term.
Figure 5: Qualitative Comparison of Texture Transfer and Style Transfer. The first four rows in the figure show the texture transfer results, while the last two rows present the style transfer results.
Geometry Preservation Branch
@@
3DGS is an explicit representation, where the scene’s ability to learn textures is closely related to the distribution of Gaussians. In low-texture regions, the density of Gaussians is typically lower. While these regions achieve good rendering quality in the original rendering process, they struggle to learn new texture appearances. The Gaussian densification strategy increases the number of Gaussians in low-texture regions during the texture migration phase, enabling the learning of complex textures. However, Gaussians are densified solely based on gradients. Since texture transfer lacks ground truth during optimization, the densification strategy may introduce erroneous Gaussians floating in space. As shown in Fig. 4, simply adding depth regularization does not solve this problem. Unlike NeRF, explicit 3D Gaussians jointly encode both geometric and color parameters. To address this issue, we propose a specialized branch for optimizing the geometric parameters of Gaussians.
Our key insight is to introduce an additional optimization objective focused on geometry preservation during training, in order to balance appearance optimization with geometric integrity. Specifically, we associate each Gaussian with an additional color parameter $c^{g}$, initialized using the original color values from the scene. During optimization, we render an image $I_{g}$ using these color parameters and optimize the Gaussian parameters by treating the content image $I_{c}$ as ground truth. The 3D Gaussian Splatting reconstruction loss function is formulated as
$\mathcal{L}_{rec}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L}_{D-SSIM},$ (11)
@@
where $\mathcal{L}_{1}$ and $\mathcal{L}_{D-SSIM}$ are calculated between the rendered image $I_{g}$ and the content image $I_{c}$. Through an optimization process with ground truth, the Gaussians in the scene are moved to their correct geometric positions.
Experiment
@@
Datasets. For the scene datasets, we utilize the LLFF dataset (mildenhall2019local) and the Tanks and Temples (T&T) dataset (knapitsch2017tanks), which are collected from the real world. Additionally, we use images from the ARF (zhang2022arf) style dataset and the DTD dataset (cimpoi2014describing) as reference image datasets.
@@
Baseline. We compare our method with the state-of-the-art 3D stylization methods, including SGSST (galerne2025sgsst), ABC-GS (liu2025abc), StyleGaussian (liu2024stylegaussian), ARF (zhang2022arf), Ref-NPR (zhang2023ref), and SNeRF (nguyen2022snerf). Specifically, SGSST, ABC-GS, and StyleGaussian are based on 3DGS, and ARF, Ref-NPR, and SNeRF are based on NeRF. StyleGaussian is a feed-forward-based method, while others are optimization-based methods.
Implementation Details
@@
We perform view-consistent color transfer (zhang2022arf) on the rendered images before and after texture transfer and use them as content images to optimize the Gaussian parameters. For the VGG (simonyan2014very) feature extractor, we employ the conv3 block of VGG-16. When associating color parameters with texture features, we set the default number of depth groups $K$ to 4, and the rotation angle $\theta$ is sampled across the full 360 degrees. The hyperparameter for AFCM is denoted as $\{\lambda_{d},\lambda_{f},\lambda_{\Phi}\}$ = {0.8, 0.8, 0.25}. For texture transfer optimization, we set $\{\lambda_{wgt},\lambda_{c},\lambda_{tv}\}$ = {2, 0.005, 0.02}. All our experiments are conducted on a single NVIDIA RTX 4090 GPU.
Qualitative Evaluation
@@
To comprehensively evaluate the effectiveness of our method, the qualitative experiments are divided into two parts: texture transfer and style transfer. The results are shown in Fig. 5.
@@
Texture Transfer. Fig. 5 shows the texture transfer results of our method compared with other methods. Visually, our results exhibit higher fidelity to the reference texture and better alignment with human visual perception. For example, in the geometrically complex orchids scene (row 1 in Fig. 5), our method produces transfer results that appear as if the texture is wrapped naturally around the surface of the orchids. StyleGaussian, as a zero-shot style transfer method, struggles to handle such complex texture transfers. The results of other optimization-based methods exhibit texture discontinuities and appear noticeably blurry, significantly deteriorating overall visual quality. This is primarily due to the inherent coupling between texture and geometry, which these methods fail to consider during optimization. In contrast, our geometry-aware approach ensures consistent texture transfer across multiple views and coherent 3D results.
Style Transfer. To validate the generalizability of our method, we selected artistic style images as reference inputs and conducted a qualitative comparison of style transfer with other methods. As shown in Fig. 5, our approach is more effective in transferring the texture elements of the style image to the 3D scene. SGSST and ABC-GS disable the Gaussian densification strategy during style transfer optimization, making it challenging to capture rich and detailed style features. For instance, in the trex scene (second-to-last row of Fig. 5), the appearance of the white wall region fails to be effectively optimized. In contrast, our method benefits from the joint effect of the AFCM and GPB. This design enables the model to preserve scene geometry while effectively learning style textures from the reference image. NeRF-based methods tend to capture more abstract stylistic elements, such as brush strokes.
Methods SSIM($\uparrow$) CLIP-score($\uparrow$) ST-LPIPS($\downarrow$) ST-RMSE($\downarrow$) LT-LPIPS($\downarrow$) LT-RMSE($\downarrow$)
Ours 0.51 0.47 0.054 0.048 0.087 0.077
SGSST 0.45 0.44 0.075 0.072 0.119 0.108
ABC-GS 0.56 0.46 0.049 0.041 0.080 0.068
StyleGaussian 0.41 0.40 0.058 0.052 0.097 0.082
ARF 0.37 0.45 0.109 0.072 0.152 0.108
Ref-NPR 0.35 0.42 0.092 0.069 0.137 0.102
SNeRF 0.48 0.36 0.075 0.057 0.127 0.090
Table 1: Quantitative Experiment on Multi-view Consistency and Content Preservation.
Quantitative Evaluation
In the 3D scene appearance editing task, maintaining multi-view consistency and preserving scene content information are crucial. To this end, we conducted extensive quantitative experiments from both aspects. We randomly selected 100 scene–reference image pairs to quantitatively evaluate our method.
Multi-view Consistency. We evaluate multi-view consistency (liu2024stylegaussian) using both short-term and long-term consistency metrics. The results are shown in Tab. 1. It can be observed that both our method and ABC-GS achieve high-quality multi-view consistency. Notably, ABC-GS disables the Gaussian densification strategy during optimization. In contrast, our method maintains multi-view consistency even after applying the densification strategy, demonstrating the effectiveness of the proposed GPB.
Content Preservation. For 3D texture transfer, it is essential to ensure that the original scene content remains recognizable while editing the scene’s appearance. We use SSIM (wang2004image) and CLIP-score (radford2021learning) to assess the preservation of content information. Specifically, SSIM evaluates the structural and informational similarity between two images at the pixel level. In contrast, CLIP-score measures the semantic similarity by computing the cosine similarity between the CLIP embeddings of the two images. As shown in Tab. 1, our proposed method achieves significantly higher scores on both evaluation metrics. In particular, our method outperforms previous approaches in terms of the CLIP-score. It indicates that incorporating geometric information enables accurate texture transfer to the scene appearance while preserving the semantic content of the scene.
Figure 6: Ablation study for GT2 Loss, AFCM, and GPB.
Ablation Study
Impact of GT2 Loss. GT2 Loss is introduced to incorporate geometric information into the optimization objective of texture transfer. As shown in Fig. 6, removing the GT2 Loss results in noticeable texture discontinuities and blurring in the texture transfer outputs. Tab. 2 further demonstrates the significance of GT2 Loss from a quantitative perspective.
Impact of AFCM. The AFCM is introduced to address the granularity mismatch between texture features and pixel-level details. It encourages texture learning in foreground and low-frequency regions, while preserving scene content in background and high-frequency areas. As illustrated in Fig. 6, in the fern scene (second row), low-texture foreground regions fail to capture style patterns without the presence of AFCM. In the truck scene (third row), which exhibits significant depth variation as a 360° environment, removing AFCM leads to noticeable degradation of geometric and appearance fidelity.
Impact of GPB. The 3D texture transfer tasks inherently lack ground truth supervision, which can introduce incorrect geometry during the scene appearance editing process. To address this, we propose GPB, which leverages the original content images to optimize the geometric parameters of Gaussians and correct inaccurate geometry. As illustrated in Fig. 6, removing GPB results in noticeable artifacts across the scene. Tab. 2 highlights the importance of GPB in maintaining content fidelity. Through GPB, our texture transfer framework achieves a balance between learning texture features and preserving scene geometry.
Ours w/o GT2 Loss w/o AFCM w/o GPB
SSIM($\uparrow$) 0.41 0.38 0.45 0.31
CLIP-score($\uparrow$) 0.39 0.36 0.38 0.37
Table 2: Quantitative Ablation Study on Content Protection. The results are obtained from 25 randomly selected experiments conducted on LLFF scenes.
Conclusion
In this paper, we introduced GT2-GS, a novel Geometry-aware Texture Transfer framework for Gaussian Splatting that achieves high-quality texture transfer results. Unlike the previous 3D stylization methods, our approach explicitly considered the intrinsic relationship between geometry and texture. We first propose the GT2 Loss, which leverages features augmented with geometric parameters and utilizes cross-view priors to guide scene optimization, enabling geometrically consistent texture transfer. AFCM addresses the granularity mismatch between features and pixels by adaptively controlling the strength of texture learning. Additionally, GPB introduces a geometry optimization objective grounded in ground-truth appearance, effectively preserving scene geometry during complex stylization. Extensive quantitative and qualitative experiments demonstrated the effectiveness of our proposed method. Moreover, our framework was capable of generating high-quality stylization results, showcasing its generalizability.
As a limitation, because optimizing rendered VGG features requires minimizing texture cosine distance while preserving content loss, which results in a texture interpolated between scene and texture geometry.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (62572194, 62472178, 62376244) and the Fundamental Research Funds for the Central Universities. This work was supported by the Shanghai Frontiers Science Center of Molecule Intelligent Syntheses. Meanwhile, we thank the financial support from Shanghai Chinafortune Co., Ltd.
ADDITIONAL ANALYSIS
User Study. Given the highly subjective nature of texture transfer evaluation, it is challenging to rely solely on objective quantitative metrics for a comprehensive assessment. Therefore, we conducted a user study to evaluate the effectiveness of the proposed method in terms of texture transfer quality. We collected 25 responses from an online questionnaire distributed through social media platforms. The questionnaire consisted of 20 groups of images along with corresponding prompts. Each group included a texture image, an original content image from a random viewpoint, and a rendered image produced by one of the methods under comparison from the same viewpoint. Each rendered image was evaluated in terms of texture alignment and visual quality, with scores ranging from 1 (lowest) to 5 (highest). As shown in Tab. 3, our method consistently outperforms the baseline approaches. Specifically, by comparing the differences across various evaluation dimensions, it is evident that users showed a clear preference for our proposed method. This indicates that our approach achieves texture transfer results that better align with human visual perception.
Runtime Cost. Runtime efficiency is crucial for 3D assets. Therefore, we evaluate the runtime cost of different methods on the LLFF dataset. As shown in Tab. 4, our method demonstrates superior efficiency in terms of both memory usage and FPS. StyleGaussian incurs higher memory costs due to the use of additional VGG feature parameters during rendering. ARF, Ref-NPR, and SNeRF are NeRF-based methods, which result in slower rendering speeds.
MORE IMPLEMENTATION DETAILS
The frequency map is derived from high-frequency energy statistics based on the Discrete Cosine Transform (DCT). Specifically, according to the scale difference between the image and the feature map, the image is divided into 8×8 blocks, and DCT is performed on each block. The frequency density of a given region is defined as the sum of the absolute DCT coefficients within the corresponding block. For the forward-facing scene dataset LLFF and the 360-degree scene dataset T&T, we adopt different experimental settings. In the color transfer stage, we perform 400 and 1000 optimization steps for LLFF and T&T scenes, respectively. During the subsequent texture transfer stage, for the LLFF dataset, we set $\{\lambda_{gt},\lambda_{c},\lambda_{tv},\lambda_{reg}\}$ = {2, 0.005, 0.02, 0.1}. For the T&T dataset, we reduce parameter $\lambda_{c}$ to 0.0005. To simplify the implementation of the geometry preservation branch, our code adopts an alternating optimization strategy between the two branches.
MORE QUALITATIVE EVALUATION
@@ Acknowledgments
More Comparison. For more qualitative evaluation, we compare our method with SGSST (galerne2025sgsst), ABC-GS (liu2025abc), StyleGaussian (liu2024stylegaussian), ARF (zhang2022arf), Ref-NPR (zhang2023ref) and SNeRF (nguyen2022snerf). Fig. 9 and Fig. 10 present the qualitative results.
Ours SGSST ABC-GS ARF
Texture Alignment($\uparrow$) 4.24 3.16 3.04 3.04
Visual Quality($\uparrow$) 3.92 3.20 3.36 3.08
Table 3: The results of user study.
Ours SGSST ABC-GS StyleGaussian ARF Ref-NPR SNeRF
Memory(GB)($\downarrow$) 1.1 1.7 1.2 6.1 1.4 1.4 1.4
FPS($\uparrow$) 151 132 133 143 8 7 7
Table 4: Runtime Cost Comparison on the LLFF Dataset.
Figure 7: Ablation Study of Content Loss.
Figure 8: Qualitative Results of Texture Orientation Control.
Figure 9: Comparison of Qualitative Results.
Figure 10: Comparison of Qualitative Results.
DISCUSSION AND LIMITATION
@@
Our method primarily consists of three key components: Geometry-aware Texture Transfer Loss (GT2 Loss), Adaptive Fine-grained Control Module (AFCM), and Geometry Preservation Branch (GPB). Among them, the proposed GT2 Loss and AFCM can be seamlessly applied to other representations, such as NeRF, TensoRF, and Plenoxel. Meanwhile, GPB provides a novel perspective for geometry-aware optimization in 3DGS-based appearance editing tasks without ground truth supervision. However, from another perspective, the upper bound of geometric accuracy after texture transfer in our framework depends on the quality of geometry obtained from the original 3DGS optimization, which may be suboptimal.
FUTURE WORK